A fighter jet takes-off from an aircraft carrier and climbs in the air with an angle of 42.1 degrees above the horizontal and a speed of 122 m/s.The speed and direction of the jet constitutes a vector quantity known as the velocity.What is the horizontal component of the plane's velocity?


1
Expert's answer
2022-04-28T13:13:19-0400

"v_x=v\sin\alpha=82~\frac ms."